\begin{thebibliography}{99}

\bibitem{Bogoliubov58}
N.N. Bogoliubov, V.V. Tolmachev, and D.V.  Shirkov,   {\it New method in the theory of
superconductivity},    (Publ. Dept. USSR Acad. of Science, Moscow, 1958); Consultants
Bureau, New York, London, Chapman and Hall, 1959, YII).

\bibitem{Morel62} P. Morel and P.W.
Anderson,  Phys.~Rev. {\bf 125}, 1263 (1962).
% Calculation of the superconducting state parameters with
%retarded electron--phonon interaction.

\bibitem{Kohn65} W. Kohn and J. M. Luttinger, Phys. Rev. Lett.
\textbf{15}, 524 (1965).

\bibitem{KaganM16} M.\ Yu. Kagan, JETP Lett. \textbf{103} 822 – 832 (2016).
% Unconventional superconductivity in low density electron systems and
%conventional superconductivity in hydrogen metallic alloys

\bibitem{Emery87} V.J. Emery,  Phys.~Rev.~Lett.  \textbf{58},
  2794 (1987).
% Theory of high-$T_c$ in oxides

\bibitem{Varma87}  C. M.~Varma, S.~Schmitt-Rink, and E.~Abrahams,
Solid State Commun.  \textbf{62}, 681 (1987).
% Charge transfer excitations and superconductivity in ionic
% metals,
\bibitem{Barabanov88} A. F. Barabanov, L.A.  Maksimov, and G. V. Uimin, JETP Lett.
 \textbf{47}, 622 (1988)]; J. Exp. Theor. Phys. \textbf{96}, 371 (1989)];
 A. F. Barabanov, R. O. Kuzian, and  L.A.  Maksimov,  J. Phys. Cond. Matter \textbf{3},
9129 (1991).
%A. F. Barabanov, L. A. Maksimov, and G.  Uimin, Pis'ma Zh. Exp. Theor. Fiz. 47,532 (1988)
% [JETP Lett. 47, 622 (1988)]; Zh. Exp. Theor. Fiz. 96, 655 (1989) [JETP 69, 371 (1989)];
%A. F. Barabanov, R. О. Kuzian, and L. А. Maksimov, J. Phys. Cond. Matter 3, 9129 (1991).

\bibitem{Valkov16} V. V. Val'kov,  D. M. Dzebisashvili,  M. M. Korovushkin,
 and A. F. Barabanov, JETP Letters \textbf{ 103},  385 (2016).
% Stability of the Superconducting  Phase in High-Tc  Superconductors with Respect
%to the Inter

\bibitem{Hayn93}  R. Hayn, V. Yushankhai,  and  S. Lovtsov,
Phys.~Rev.~B {\bf 47}, 5253 (1993).
%Analysis of the singlet-triplet  model for the copper
%oxide plane within the paramagnetic state.

\bibitem{Belinicher94}   V. I. Belinicher and A. L. Chernyshev, Phys. Rev. B {\bf 49},
 9746 (1994).
%Consistent low-energy reduction of the three-band model for copper oxides with O-O
%hopping to the effective t-J model

\bibitem {Feiner96} L. F. Feiner, J. H. Jefferson and R. Raimondi,
  Phys. Rev. B  {\bf 53}, 8751 (1996)
% Effective single-band models for  the high-$T_c$ cuprates.  I. Coulomb
%interactions.

\bibitem{Yushankhai97} V. Yushankhai, V.S. Oudovenko,
and R. Hayn,   Phys.~Rev.~B {\bf 55}, 15562 (1997).
% Proper reduction scheme to  an extended $t$-$J$ model
%and the hole dispersion in  Sr$_2$CuO$_2$Cl$_2$.

\bibitem{Korshunov05}  M. M. Korshunov, V. A. Gavrichkov,  S. G. Ovchinnikov, I. A. Nekrasov,
Z. V. Pchelkina, and V. I. Anisimov,  Phys.~Rev.~B {\bf  72},  165104 (2005).
% Hybrid LDA and generalized tight-binding method for electronic structure calculations of strongly
%correlated electron systems

\bibitem{Hubbard65} J. Hubbard,
Proc.~Roy. Soc. A (London) \textbf{285}, 542 (1965).
% Electron correlations in narrow energy bands.  IV.  The atomic
% representation.

\bibitem{Plakida13}   N.M.~Plakida  and  V.S.~Oudovenko,
Eur. Phys. J. B \textbf{86}, 115 (2013).
%On the theory of superconductivity in the
%extended Hubbard model: Spin-fluctuation pairing

\bibitem{Plakida14}  N.M. Plakida and V.S. Oudovenko,   J. Exp. Theor. Phys.
{\bf 146}, 631  (2014).
%Kinematic spin-fluctuation mechanism of high-temperature superconductivity

\bibitem{Zaitsev87} R.O. Zaitsev,  and  V.A. Ivanov, JETP Lett.
{\bf 46} , S116 (1987), Soviet Phys. Solid State \textbf{29},    2554  (1987),
% On the possibility of pair condensation in the Hubbard model.
 Ibid. {\bf 29},   3111 (1987), Int. J. Mod. Phys. B \textbf{5},  153  (1988).
% Superconductivity in the Hubbard model.

\bibitem{Zaitsev04} R. O. Zaitsev, J. Exp. Theor. Phys. \textbf{98}, 780 (2004).
%Peculiarities of the Electron Mechanism of Superconductivity

\bibitem{Valkov11}  V. V. Val'kov and M. M. Korovushkin, J. Exp. Theor.
Phys. \textbf{112}, 108 (2011).
% Splitting of the Lower Subband of Hubbard Fermions in the
%Shubin–Vonsowsky Model under the Influence of Strong Intersite Correlations

\bibitem{Plakida89} N.M. Plakida, V.Yu. Yushankhai, and
I.V. Stasyuk,  Physica C  \textbf{160}, 80 (1989).
% On the role of kinematical and exchange interactions in
%superconducting pairing of electrons in the Hubbard model,

\bibitem{Yushankhai91} V.Yu. Yushankhai, N.M. Plakida,  and P. Kalinay,
 Physica C \textbf{174}, 401 (1991).
% Superconducting pairing in the mean-field
%approximation for the $t-J$ model: Numerical analysis.

\bibitem{Valkov03} V. V. Val'kov,  D. M. Dzebisashvili, and A.S. Kravtsov,
JETP Letters \textbf{77},  505 (2003);  Doklady  USSR Acad. of Science
 {\bf 393}, 1 (2003).
% Spectral Representations and Description of a Superconducting State with
% S-Type Order Parameter  $\Delta(k)$.

\bibitem{Kubo57}  R. Kubo, J. Phys. Soc. Jpn. {\bf 12}, 570 (1957).

\bibitem{Suzuki71} M. Suzuki, Physica {\bf 51}, 277 (1971).

\bibitem{Huber77}  D.L. Huber, Physica A, {\bf 87 A}, 199 (1977).


\bibitem{Mancini03} F. Mancini, A. Avella, Eur. Phys. J. B  {\bf 36}, 37 (2003).
% Equation of motion method for composite field operators

\bibitem{Avella07} A. Avella, F. Mancini, J. Phys.: Condens. Matter {\bf 19}, 255209 (2007)
%The 2D Hubbard model and the pseudogap: a COM(SCBA) study



\end{thebibliography}

\documentclass[aps,twocolumn,showpacs,floatfix,amssymb]{revtex4}
\usepackage{graphicx}
\begin{document}

\title{On the  Coulomb interaction in\\
superconducting pairing in cuprates }
\author{Nikolay M. Plakida }
 \affiliation{Joint Institute for Nuclear Research,
 141980 Dubna, Russia}

\date{\today}

\begin{abstract}

We discuss a reduction of   superconducting transition temperature by the intersite Coulomb repulsion for the  $d$-wave pairing  in cuprates.   We compare the results found for  the spin-fermion model and the
extended Hubbard model. We argue that in both the models the  $d$-wave superconducting
transition temperature is reduced by the Coulomb repulsion of holes in different unit
cells. We also show that in the strong correlation limit the  $s$-wave superconductivity
cannot occur due to the kinematic restriction  of no double occupancy in the Hubbard
subbands.
\end{abstract}

\pacs{74.20.Mn, 71.27.+a, 71.10.Fd, 74.72.-h}

 \maketitle

It is commonly believed that the Coulomb interaction  is detrimental to
superconductivity. In particular, for low-temperature superconductors with the $s$-wave
pairing  mediated by electron-phonon coupling the retardation effect renormalizes the
Coulomb interaction which makes it feasible to obtain a finite superconducting
$T_c$~\cite{Bogoliubov58,Morel62}. For  electronic pairing mechanisms,  the retardation
effect is ineffective and the Coulomb interaction suppresses the $s$-wave pairing. Only
superconducting pairing  with higher orbital momenta,  $p, d, f, \ldots $, can occur in
the Fermi-liquid as originally was proposed by Kohn and Luttinger~\cite{Kohn65} (for a
review see \cite{KaganM16}).

The cuprate superconductors are the Mott-Hubbard (more accurately, charge-transfer) doped
insulators caused by the large Coulomb interaction $U_d$ on copper sites.  In this case
the interaction  $U_d$ should be taken into account rigorously  in considering the
electronic structure of cuprates.  The most frequently  used  is the three-band $p$--$d$
model for  Cu$\, 3\, d(x^{2}-y^{2})$-states and O$2p_{\sigma} (x,y)$-states in the
CuO$_2$ plane~\cite{Emery87,Varma87}.  To obtain a tractable model for description of
low-energy electronic excitations the  $p$--$d$ model can be reduced to  simpler  models.

In particular,   in the spin-fermion model (SFM) the high-energy excitations on copper
sites are excluded, which results in a conduction band  for oxygen holes on the O$2p_{\sigma}
(x,y)$ orbitals interacting with localized copper spins $S= 1/2$ in the CuO$_2$ plane
(see, e.g.,~\cite{Barabanov88}). In Ref.~\cite{Valkov16}, the SFM was used to consider the
$d$-wave   superconducting pairing for spin-polarons. It was found that the Coulomb
interaction $V_{pp}$ between holes on the nearest neighbor oxygen sites gives no
contribution to the $d$-wave pairing by  symmetry reason.  This was considered as a proof
of stability of the $d$-wave pairing towards  the intersite Coulomb repulsion. However,
the authors have neglected the Coulomb interaction between  holes in  different unit
cells,  which conventionally reduces  the superconducting $T_c$.

Another approach is   based on the cell-cluster perturbation theory (see
\cite{Hayn93,Belinicher94,Feiner96,Yushankhai97,Korshunov05}).  In the theory,  the
spectrum of  electronic excitations in the unit cell  CuO$_4$ is rigorously calculated by
an  exact diagonalization of the copper and oxygen energy states taking into account all
relevant  Coulomb interactions, $U_d,\, U_p,\, V_{pd}, \, V_{pp}$, and hybridizations,
$t_{pd}, \, t_{pp}$. Considering the lowest energy states close to the Fermi level, the
singly occupied $d (x^2 - y^2)$ states  and doubly occupied singlet $p$-$d$ hole states,
the extended Hubbard model (EHM) can be formulated  for two  Hubbard subbands with the
hopping parameter between different unit cells   $ t \sim 0.3 t_{pd}$ and the intersite
Coulomb repulsion $V \sim 0.5\, t $~\cite{Feiner96}.

In the limit of strong correlations the projected (Hubbard) electronic operators should
be used~\cite{Hubbard65}. They have nonfermionic commutation relations, as e.g., the commutation relation for the Hubbard operators in the singly occupied subband, $ X_i^{0 \sigma } = a_{i\sigma}(1-
N_{i\bar\sigma})$ where $N_{i \sigma} = a^\dag_{i\sigma} a_{i\sigma}, \; \sigma = \pm 1,\, \bar\sigma = -
\sigma $, reads
\begin{equation}
X_i^{0 \sigma }   X_j^{\sigma0} +  X_j^{\sigma0}X_i^{0 \sigma }  = \delta_{ij}(1 - N_{i
\sigma} /2 + \sigma S_i^z).
   \label{1a}
\end{equation}
This results in the kinematic interaction for electrons,  which is determined by electron
scattering on charge (number $N_{i \sigma}$) and spin $S_i^\alpha$ fluctuations with the
coupling of an order of the hopping parameter $\, t$.

In Refs.~\cite{Plakida13,Plakida14}, the EHM   was studied within the strong coupling
superconducting theory. It was shown that the spin-fluctuation  pairing induced by the
kinematical interaction in the second order of $\, t \,$ results in the $d$-wave
superconductivity with high-$T_c$. In Ref.~\cite{Plakida13}, it was found  that the
intersite Coulomb repulsion $V \sim 0.5\, t $ in cuprates is not strong enough to
suppress  the $d$-wave superconductivity.  To prove this pairing mechanism, in
Ref.~\cite{Plakida14} we consider a much stronger than in cuprates Coulomb interaction
$V$. Only for $V$  larger than the coupling constant for the spin-fluctuation pairing, $V  \gtrsim 4\,t$,  the $d$-wave pairing can be fully suppressed.

Now we comment on  the $s$-wave superconducting pairing in  cuprates.    The $s$-wave
pairing induced by the kinematical interaction in the limit of strong correlations was
originally proposed  in Refs.~\cite{Zaitsev87}.  In Refs.~\cite{Zaitsev04, Valkov11},  the superconducting pairing in the two-dimensional Hubbard model  was studied. It
was found that this pairing is robust in respect to the  intersite  Coulomb interaction
$V$. However, in these papers only the first order of the kinematical interaction $\propto t $ was
considered, which results in the  $s$-wave  pairing with the symmetric superconducting gap
 $ \,\Delta_{\sigma}(q_x, q_y) =  \Delta_{\sigma}(q_y,
q_x)\,$. However, in this case the well known constraint of ``no double occupancy'' in
strongly correlated systems is violated. First, it was pointed out in
Refs.~\cite{Plakida89,Yushankhai91} for the $t$-$J$ model and then in
Ref.~\cite{Plakida14} for the EHM model. This constraint can be  formulated in terms of a
specific  relation for the anomalous (pair) correlation function for the Hubbard
operators. It is easy to verify that a product of two Hubbard operators for the singly
occupied subband equals zero: $\, X_i^{0 \sigma }  X_i^{0 \bar{\sigma}} =
a_{i\sigma}(1- N_{i\bar\sigma})\, a_{i\bar\sigma}(1- N_{i\sigma}) = 0$. Therefore,   the
corresponding single-site pair correlation function should vanish:
\begin{equation}
F_{ii,\sigma} =
  \langle  X_i^{0 \sigma }   X_i^{0 \bar{\sigma}} \rangle =
    \frac{1}{N} \sum_{{\bf q}}\, \langle  X_{\bf q}^{0 \sigma }
    X_{-{\bf q}}^{0 \bar{\sigma}} \rangle  \equiv 0.
\label{1}
\end{equation}
The symmetry of the Fourier-component of the  pair correlation function
$\,F_{\sigma}({\bf q}) = \langle  X_{\bf q}^{0 \sigma } X_{-{\bf q}}^{0 \bar{\sigma}}
\rangle  \,$ has the symmetry of the superconducting order parameter, i.e., the gap
function. For the tetragonal lattice for the $d$-wave pairing $ \,F_{\sigma}(q_x, q_y) =
- F_{\sigma}(q_y, q_x)\,$ and the condition (\ref{1}) after integration over $q_x, q_y$
is fulfilled. For the $s$-wave pairing $ \,F_{\sigma}(q_x, q_y) =  F_{\sigma}(q_y,
q_x)\,$ and the condition (\ref{1}) is violated. The same condition holds for the pair
correlation function for the second Hubbard subband, $\langle  X_i^{\sigma 2}
X_i^{\bar{\sigma}2} \rangle = 0$. Therefore, the $s$-wave pairing in both the Hubbard
subbands is prohibited in the limit of strong correlations.

To overcome the restriction (\ref{1}) in Refs.~\cite{Valkov03}  it was proposed  to
consider the modified time-dependent pair correlation function:
\begin{equation}
\tilde{F}_{ii,\sigma} (t) =
  \int_{- \infty}^{+\infty} d \omega \, {\rm e}^{i\omega t}\,
  \tilde{J}_{ii,\sigma}(\omega) ,
\label{2}
\end{equation}
where
\begin{equation}
  \tilde{J}_{ii,\sigma}(\omega) = {J}_{ii,\sigma}(\omega)-
 \delta(\omega) \;\int_{- \infty}^{+\infty} d \omega_1 \,  J_{ii,\sigma}(\omega_1)  .
\label{3}
\end{equation}
The  spectral density $\,{J}_{ii,\sigma}(\omega)\,$ determines the original correlation
function
\begin{equation}
{F}_{ii,\sigma} (t) =
  \langle  {X_{i}^{0\sigma}}(t)  {X_{i}^{0 \bar{\sigma}}} \rangle =
  \int_{- \infty}^{+\infty} d \omega \, {\rm e}^{i\omega t}\,
 {J}_{ii,\sigma}(\omega) .
\label{4}
\end{equation}
For the modified spectral density  (\ref{3}) the condition  (\ref{1}) is trivially
satisfied for any spectral function $\, J_{ii,\sigma}(\omega) \,$ and the restriction on
the $s$-wave pairing seems to be  lifted. However, thy spectral density (\ref{3}) results
in the nonergodic behavior~\cite{Kubo57}  of the pair correlation function (\ref{2}):
\begin{eqnarray}
C_{ii,\sigma} =  \lim_{t \to \infty} \tilde{F}_{ii,\sigma} (t)
 = - \frac{1}{N} \sum_{{\bf q}}\, \langle X_{\bf q}^{0 \sigma }
    X_{-{\bf q}}^{0 \bar{\sigma}} \rangle ,
\label{5}
\end{eqnarray}
where the conventional pair correlation function decays in the limit ${t \to \infty} $
due to finite life-time effects
\begin{equation}
 \lim_{t \to \infty} {F}_{ii,\sigma} (t)  =
\int_{- \infty}^{+\infty} d \omega \, {\rm e}^{i\omega t}\, J_{ii,\sigma}(\omega) = 0.
\label{6}
\end{equation}
The nonergodic behavior of the modified pair correlation function (\ref{2})  contradicts 
the basic  properties of physical systems and appears for some pathological models with
local integrals of motion~\cite{Suzuki71,Huber77}. In that case, the nonergodic constants
can be found from $1/\omega$ poles of the anticommutator or causal Green functions, as described for the Hubbard model for spin or charge excitations in Refs.~\cite{Mancini03,Avella07} contrary to the arbitrary definition (\ref{3}). Therefore, the statement given in Refs.~\cite{Valkov03}: ``The inclusion of a singular contribution to the spectral intensity of the anomalous correlation function regains the sum rule and remove the unjustified forbidding of the $s$-symmetry order parameter in superconductors with
strong correlations''  cannot be accepted.\\

To conclude, the contradiction between the theoretical and experimental results claimed
in Ref.~\cite{Valkov16}: a more stable  $s$-wave superconducting  pairing with respect to
the intersite Coulomb interaction found in  Refs.~\cite{Zaitsev04, Valkov11}, and a
strong suppression of  the $d$-wave pairing found in Refs.~\cite{Plakida13,Plakida14}, in
fact, is absent. The $s$-wave pairing in the limit of strong correlations is prohibited
due to the kinematical restriction (\ref{1}), while the $d$-wave pairing found within the
EHM can be suppressed only for unphysically large Coulomb interaction $V$, as shown in
Refs.~\cite{Plakida14}. The cancellation of the intersite Coulomb interaction for the
$d$-wave pairing in SFM was found in Ref.~\cite{Valkov16} only  for the nearest-neighbor
oxygen sites, which does not prevent suppression of $d$-wave pairing
due to the  Coulomb interaction for  oxygen sites in different unit cells as in EHM.\\

The author thanks V. Yushankhai for useful discussions.

\end{document}